OpenStudy (anonymous):

A number exceeds four times it reciprocal by 3. Find the number.

OpenStudy (henryblah):

Let number be x. Reciprocal = 1/x. x=4*(1/x) + 3 Solve.

OpenStudy (mertsj):

Let x = the number Then 1/x= the reciprocal. And: \[x=4(\frac{1}{x})+3\] Solve that equation

OpenStudy (henryblah):

XD Similar stuff :)